# Exact and Linear Convergence for Federated Learning under Arbitrary Client Participation is Attainable

### v2

@@ C.2 Convergence Proof
[…]
@@
The subsequent proof follows a standard framework for analyzing decentralized algorithms. It initially establishes a descent lemma, demonstrating that the virtual weighted iterates $\bar{x}_{k}$ progressively approach a neighborhood of the optimal solution in each iteration. Subsequently, a consensus lemma is established, showing that the individual client iterates, $x_{k,i}$, gradually converge towards this weighted iterate $\bar{x}_{k}$. Finally, by combining these two lemmas, we will derive the overall convergence theorem.
C.2.1 Descent Lemma
@@
where we discarded the $(6\eta^{2}L-2\eta)\big{(}F(\bar{x}_{k})-F(x^{\star})\big{)}$ term since it is always negative. $\hfill\qed$
C.2.2 Consensus Lemma
@@
$\left\|q^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{k^{\prime}})-\nabla f_{i}(x_{k^{\prime},i})\right\|^{2}$
@@
$\hfill\qed$
C.2.3 Proof of Convergence Theorem 2
@@
where we introduce the conditional number $\kappa=L/\mu$. $\hfill\qed$
@@
Appendix D Proof of the Convergence of FOCUS
D.1 Reformulate the Recursion
@@
Iter. $k$ 0 1 2 $\cdots$ $\tau$ $\tau+1$cdots $r\tau-1$ $r\tau$ $r\tau+1$cdots
$R_{k}$ Init. $R(S_{1})$ $I$cdots $I$ $R(S_{2})$cdots $I$ $R(S_{r})$ $I$cdots
$C_{k}$ Init. $I$ $I$cdots $C(S_{1})$ $I$cdots $C(S_{r-1})$ $I$ $I$cdots
@@
Iter. $k$1 0 1 2 $\cdots$ $\tau$ $\tau+1$cdots $r\tau-1$ $r\tau$ $r\tau+1$cdots
$C_{k}$ Init. $I$ $I$cdots $C(S_{1})$ $I$cdots $I$ $C(S_{r-1})$ $I$cdots
$R_{k}$ Init. $R(S_{1})$ $I$ $I$cdots $R(S_{2})$ $I$cdots $I$ $R(S_{r})$ $I$cdots
@@ D.2 Useful Observations
$\bar{x}_{k+1}=\begin{cases}\bar{x}_{k}&\;\;k\neq r\tau\\ \bar{x}_{k}-\eta\bar{y}_{k+1}&\;\;k=r\tau\\ \end{cases},\;\;\;\;\;\bar{y}_{k+1}=\begin{cases}\bar{y}_{k}&\;\;k\neq r\tau\\ \mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{k})&\;\;k=r\tau\\ \end{cases}$ (71)
@@
D.3 Descent Lemma for FOCUS
@@
[…]
@@
[…]
@@
[…]
@@
where the first equation means the difference between the model before pulling and the server’s model, and the second equation means the difference between the clients’ models after pulling and the server’s model. The weighted $Q$ and $I-Q$ (ref. (73)-(74)) are strictly smaller than 1, which can make the convergence proof tighter. But for simplicity, we just loosen it to 1 and got
@@
where the last inequality holds when $\eta\leq\frac{1}{3(\tau-1)L}$. $\hfill\qed$
D.4 Consensus Lemma for FOCUS
@@
$\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
$\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
$\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
$\hfill\qed$
D.5 Proof of the Convergence of FOCUS ($\mu-$Strong Convexity)
@@
where (102) holds when $\eta\leq\frac{2q_{\min}}{3N\left(\frac{8L^{2}}{\mu}+\frac{\mu}{2}\right)}$. $\hfill\qed$
@@
D.6 Proof of the Convergence of FOCUS (Non-Convexity with PL Assumption)
@@
where the second inequality follows $\eta\leq\min\{\frac{1}{4LN},\frac{1}{2(\tau-1)LN}\}$ and the last inequality we split $-\eta N\|\nabla F(\bar{x}_{r\tau+1})\|^{2}$ into two parts and applying PL-condition for one part. Next, we minus $F^{\star}$ on the both sides to obtain:
$F(\bar{x}_{(r+1)\tau+1})-F^{\star}$
@@
Therefore, we conclude that FOCUS achieves the linear convergence rate of $(1-\eta\beta N)$ in the nonconvex case with PL condition. $\hfill\qed$
D.7 Proof of the Convergence of FOCUS (General Non-Convexity)
@@
$-\eta\left\langle\nabla F(\bar{x}_{r\tau+1}),\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\right\rangle$
@@
$\frac{\eta N}{2}(1-2\eta NL)\mathbb{E}\hskip 1.42271pt\|\nabla F(\bar{x}_{r\tau+1})\|^{2}$
@@
$\frac{\eta N}{2}(1-2\eta NL)\mathbb{E}\hskip 1.42271pt\|\nabla F(\bar{x}_{r\tau+1})\|^{2}$
@@
Taking the summation from $r=0$ to $R-1$, we have
@@
[…]
where the last inequality follows $\eta\leq\frac{q_{\min}}{16L\sqrt{2N}}$. $\hfill\qed$
Appendix E Extensions to Stochastic Gradient Case ( SG-FOCUS )
@@
5:        $x_{0,i}^{(r)}\Leftarrow x_{r},\;y_{0,i}^{(r)}\Leftarrow 0$triangleright Pull model $x_{r}$ from server while $y_{r}$ is NOT pulled
@@
7:           $y_{t+1,i}^{(r)}\!=\!y_{t,i}^{(r)}\!+\!\nabla f_{i}(x_{t,i}^{(r)};\xi_{t})\!-\!\nabla f_{i}(x_{t-1,i}^{(r)};\xi_{t-1})$triangleright Current and last stochastic grad.
@@
11:     $y_{r+1}=y_{r}+\sum_{i\in S_{r}}y^{(r)}_{\tau,i}$triangleright Pushing $y^{(r)}_{i,\tau}$ to server (Not averaging)
12:     $x_{r+1}=x_{r}-\eta y_{r+1}$triangleright Client model $x^{(r)}_{\tau,i}$ is NEVER pushed.
@@
E.1 Descent Lemma for SG-FOCUS
@@
[…]
@@
where the last inequality follows $\eta\leq\min\{\frac{1}{\mu},\frac{\mu}{128(\tau-1)^{2}L^{2}N},\frac{1}{2\sqrt{2}(\tau-1)L}\}=\min\{\frac{\mu}{128(\tau-1)^{2}L^{2}N},\frac{1}{2\sqrt{2}(\tau-1)L}\}$. $\hfill\qed$
E.2 Consensus Lemma for SG-FOCUS
@@
$\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
Putting (E.2), (131), (132) back to (129) and selecting $\rho=\frac{1-q_{\min}}{1-q_{\min}/2}=1-\frac{q_{\min}}{2-q_{\min}}$, we obtain
$\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
$\left(1-\frac{q^{2}_{\min}}{4}\right)\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
$\hfill\qed$
E.3 Proof of the Convergence of SG-FOCUS ($\mu-$Strong Convexity)

### v3

@@ D.2 Convergence Proof
$\mathbb{E}\hskip 1.42271pt\|\bar{x}_{k+1}-x^{\star}\|^{2}=$ $\mathbb{E}\hskip 1.42271pt\|\bar{x}_{k+1}-\mathbb{E}\hskip 1.42271pt\bar{x}_{k+1}+\mathbb{E}\hskip 1.42271pt\bar{x}_{k+1}-x^{\star}\|^{2}$
$=$ $\mathbb{E}\hskip 1.42271pt\|\bar{x}_{k+1}-\mathbb{E}\hskip 1.42271pt\bar{x}_{k+1}\|^{2}+\|\mathbb{E}\hskip 1.42271pt\bar{x}_{k+1}-x^{\star}\|^{2}$
$=$ $\left\{\begin{aligned} &\left\|\bar{x}_{k}-\eta q^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{k})-x^{\star}\right\|^{2},&k\neq r\tau\\ &\mathbb{E}\hskip 1.42271pt\left\|\bar{x}_{k+1}-q^{\mathsf{T}}{\boldsymbol{y}}_{k+1}\right\|^{2}+\left\|\bar{x}_{k}-\eta q^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{k})-x^{\star}\right\|^{2},&k=r\tau,\end{aligned}\right$. (48)
@@
The subsequent proof follows a standard framework for analyzing decentralized algorithms. It initially establishes a descent lemma, showing that the virtual weighted iterates $\bar{x}_{k}$ progressively approach a neighborhood of the optimal solution in each iteration. Subsequently, a consensus lemma is established, showing that the individual client iterates, $x_{k,i}$, gradually converge towards this weighted iterate $\bar{x}_{k}$. Finally, by combining these two lemmas, we will derive the overall convergence theorem.
D.2.1 Descent Lemma
@@
$\hskip-14.22636pt\left\|\bar{x}_{k}-\eta q^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{k})-x^{\star}\right\|^{2}$
@@
[…]
D.2.2 Consensus Lemma
@@
$\hskip-14.22636pt\left\|q^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{k^{\prime}})-\nabla f_{i}(x_{k^{\prime},i})\right\|^{2}$
@@
[…]
D.2.3 Proof of Convergence Theorem 3
@@
As we discussed in the main context, these three terms have their own meanings. We can easily establish the following corollaries. Note $\delta_{q}^{2}=0$ under the uniform sampling case. We have
@@
Appendix E Proof of the Convergence of FOCUS
E.1 Reformulate the Recursion
@@
Iter. $k$ 0 1 2 $\cdots$ $\tau$ $\tau+1$ $\cdots$ $r\tau-1$ $r\tau$ $r\tau+1$ $\cdots$
$R_{k}$ Init. $R(S_{1})$ II $\cdots$ II $R(S_{2})$ $\cdots$ II $R(S_{r})$ II $\cdots$
$C_{k}$ Init. II II $\cdots$ $C(S_{1})$ II $\cdots$ $C(S_{r-1})$ II II $\cdots$
@@
Iter. $k$ -1 0 1 2 $\cdots$ $\tau$ $\tau+1$ $\cdots$ $r\tau-1$ $r\tau$ $r\tau+1$ $\cdots$
$C_{k}$ - Init. II II $\cdots$ $C(S_{1})$ II $\cdots$ II $C(S_{r-1})$ II $\cdots$
$R_{k}$ Init. $R(S_{1})$ II II $\cdots$ $R(S_{2})$ II $\cdots$ II $R(S_{r})$ II $\cdots$
@@ E.2 Useful Observations
$\bar{x}_{k+1}=\begin{cases}\bar{x}_{k}&\;\;k\neq r\tau\\ \bar{x}_{k}-\eta\bar{y}_{k+1}&\;\;k=r\tau\\ \end{cases},\;\;\;\;\;\bar{y}_{k+1}=\begin{cases}\bar{y}_{k}&\;\;k\neq r\tau\\ \mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{k})&\;\;k=r\tau\\ \end{cases},$ (79)
@@
E.3 Descent Lemma for FOCUS
@@
$=$ $\|\bar{x}_{r\tau+1}-x^{\star}\|^{2}-2\eta\mathbb{E}\hskip 1.42271pt\big\langle\bar{x}_{r\tau+1}-x^{\star},\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\big\rangle+\eta^{2}\mathbb{E}\hskip 1.42271pt\left\|\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\right\|^{2}$
$\leq$ $\|\bar{x}_{r\tau+1}-x^{\star}\|^{2}-2\eta\mathbb{E}\hskip 1.42271pt\big\langle\bar{x}_{r\tau+1}-x^{\star},\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\big\rangle+2\eta^{2}\mathbb{E}\hskip 1.42271pt\left\|\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})-\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}(\mathds{1}\bar{x}_{r\tau+1})\right\|^{2}$
@@
$\leq$ $\|\bar{x}_{r\tau+1}-x^{\star}\|^{2}-2\eta\mathbb{E}\hskip 1.42271pt\big\langle\bar{x}_{r\tau+1}-x^{\star},\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\big\rangle+2\eta^{2}\mathbb{E}\hskip 1.42271pt\left\|\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})-\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}(\mathds{1}\bar{x}_{r\tau+1})\right\|^{2}$
@@
$\;\;\;+2\eta\left(\frac{1}{\mu}+\eta\right)\mathbb{E}\hskip 1.42271pt\left\|\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})-\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}(\mathds{1}\bar{x}_{r\tau+1})\right\|^{2}$ (90)
@@
$\hskip-14.22636pt\mathbb{E}\hskip 1.42271pt\left\|\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}(\mathds{1}\bar{x}_{r\tau+1})-\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\right\|^{2}$
@@
where the first equation means the difference between the model before pulling and the server’s model, and the second equation means the difference between the clients’ models after pulling and the server’s model. The weighted $Q$ and IQI-Q (ref. (81)-(82)) are strictly smaller than 1, which can make the convergence proof tighter. But for simplicity, we just loosen it to 1 and got
@@
[…]
E.4 Consensus Lemma for FOCUS
@@
$\hskip-14.22636pt\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
$\hskip-14.22636pt\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
$\hskip-5.69054pt\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
[…]
E.5 Proof of the Convergence of FOCUS ($\mu$-Strong Convexity)
@@
[…]
@@
E.6 Proof of the Convergence of FOCUS (Non-Convexity with PL Assumption)
@@
where the second inequality follows $\eta\leq\min\{\frac{1}{4LN},\frac{1}{2(\tau-1)LN}\}$ and the last inequality we split $\eta N\|\nabla F(\bar{x}_{r\tau+1})\|^{2}$ into two parts and applying PL-condition for one part. Next, we minus $F^{\star}$ on the both sides to obtain:
$\hskip-14.22636ptF(\bar{x}_{(r+1)\tau+1})-F^{\star}$
@@
[…]
E.7 Proof of the Convergence of FOCUS (General Non-Convexity)
@@
$\hskip-14.22636pt-\eta\left\langle\nabla F(\bar{x}_{r\tau+1}),\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\right\rangle$
@@
$\hskip-14.22636pt\frac{\eta N}{2}(1-2\eta NL)\mathbb{E}\hskip 1.42271pt\|\nabla F(\bar{x}_{r\tau+1})\|^{2}$
@@
$\hskip-14.22636pt\frac{\eta N}{2}(1-2\eta NL)\mathbb{E}\hskip 1.42271pt\|\nabla F(\bar{x}_{r\tau+1})\|^{2}$
@@
Taking the summation from $r=0$ to R1R-1, we have
@@
$\Longrightarrow\;\;\;\;\;\;$ $\left(1-\frac{256\eta^{2}L^{2}N}{q_{\min}^{2}}\right)\frac{1}{R}\sum_{r=0}^{R-1}A_{r}\leq\frac{4}{\eta N}(B_{0}-B_{R})$
$\Longrightarrow\;\;\;\;\;\;$ $\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\hskip 1.42271pt\|\nabla F(\bar{x}_{r\tau+1})\|^{2}\leq\frac{8\Big(F(\bar{x}_{1})-F(\bar{x}^{\star})\Big)}{\eta NR},$ (129)
@@
Appendix F Extensions to Stochastic Gradient Case ( SG-FOCUS )
@@
5:   $x_{0,i}^{(r)}\Leftarrow x_{r},\;y_{0,i}^{(r)}\Leftarrow 0$ $\triangleright$ Pull model $x_{r}$ from server while $y_{r}$ is NOT pulled
@@
7:    $y_{t+1,i}^{(r)}\!=\!y_{t,i}^{(r)}\!+\!\nabla f_{i}(x_{t,i}^{(r)};\xi_{t})\!-\!\nabla f_{i}(x_{t-1,i}^{(r)};\xi_{t-1})$ $\triangleright$ Current and last stochastic grad.
@@
11:  $y_{r+1}=y_{r}+\sum_{i\in S_{r}}y^{(r)}_{\tau,i} \triangleright$ Pushing $y^{(r)}_{i,\tau}$ to server (Not averaging)
12:  $x_{r+1}=x_{r}-\eta y_{r+1}$ $\triangleright$ Client model $x^{(r)}_{\tau,i}$ is NEVER pushed.
@@
F.1 Descent Lemma for SG-FOCUS
@@
$=$ $\|\bar{x}_{r\tau+1}-x^{\star}\|^{2}-2\eta\mathbb{E}\hskip 1.42271pt\big\langle\bar{x}_{r\tau+1}-x^{\star},\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau};\xi_{(r+1)\tau})\big\rangle+\eta^{2}\mathbb{E}\hskip 1.42271pt\left\|\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau};\xi_{(r+1)\tau})\right\|^{2}$
@@
$\hskip-14.22636pt\mathbb{E}\hskip 1.42271pt\left\|\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}(\mathds{1}\bar{x}_{r\tau+1})-\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\right\|^{2}$
@@
$\hskip-14.22636pt\mathbb{E}\hskip 1.42271pt\left\|\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}(\mathds{1}\bar{x}_{r\tau+1})-\mathds{1}^{\mathsf{T}}\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{(r+1)\tau})\right\|^{2}$
@@
where the last inequality follows:
$\eta\leq\min\{\frac{1}{\mu},\frac{\mu}{128(\tau-1)^{2}L^{2}N},\frac{1}{2\sqrt{2}(\tau-1)L}\}=\min\{\frac{\mu}{128(\tau-1)^{2}L^{2}N},\frac{1}{2\sqrt{2}(\tau-1)L}\}.$
[…]
F.2 Consensus Lemma for SG-FOCUS
@@
$\hskip-14.22636pt\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
Putting (F.2), (139), (140) back to (137) and selecting $\rho=\frac{1-q_{\min}}{1-q_{\min}/2}=1-\frac{q_{\min}}{2-q_{\min}}$, we obtain
$\hskip-14.22636pt\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
$\hskip-14.22636pt\left(1-\frac{q^{2}_{\min}}{4}\right)\mathbb{E}\hskip 1.42271pt\|\mathds{1}\bar{x}_{r\tau+1}-{\boldsymbol{x}}_{r\tau}\|^{2}_{F}$
@@
[…]
F.3 Proof of the Convergence of SG-FOCUS ($\mu$-Strong Convexity)